OpenStudy (anonymous):

What is the solution for 7-9b=-8b+8

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

solve for b

OpenStudy (anonymous):

-9b+8b=8-7 -1b=1 b=1/-1 b=-1